OpenStudy (anonymous):

The slope of the line tangent to the curve y^2 + (xy+1)^3 = 0 at (2, -1) is

OpenStudy (amistre64):

the implicit derivative

OpenStudy (anonymous):

is that when you take the derivative and include y' ?

OpenStudy (amistre64):

yes, which afterall is the same as taking a regular derivative, only "y" isnt easy to pull out

OpenStudy (anonymous):

right. That's where I went wrong. I tried to take the derivative, but I must have messed up somewhere because I got a really odd number that's not one of my answer choices.

OpenStudy (anonymous):

once I get the derivative, can I plug in immediately for x and y, or do I have to reduce / combine?

OpenStudy (amistre64):

\[\frac{d}{dx}(y^2 + (xy+1)^3 = 0)\] \[\frac{d}{dx}y^2 + \frac{d}{dx}(xy+1)^3 = \frac{d}{dx}0\] \[\frac{dy}{dx}2y + \frac{d}{dx}(xy+1)\ *3(xy+1)^2 = 0\] \[\frac{dy}{dx}2y + \left(\frac{d}{dx}(xy)+\frac{d}{dx}1\right)\ * 3(xy+1)^2 = 0\] \[\frac{dy}{dx}2y + \left(\frac{d}{dx}xy+x\frac{d}{dx}y+\frac{dx}{dx}\right)\ * 3(xy+1)^2 = 0\] \[\frac{dy}{dx}2y + \left(\frac{dx}{dx}y+x\frac{dy}{dx}+\frac{dx}{dx}0\right)\ * 3(xy+1)^2 = 0\] \[\frac{dy}{dx}2y + \left(y+x\frac{dy}{dx}\right)\ * 3(xy+1)^2 = 0\] \[\frac{dy}{dx}2y + \left(3y+3x\frac{dy}{dx}\right) (xy+1)^2 = 0\] this gets messy lol

OpenStudy (amistre64):

once you get the derivate, use the point to calibrate it, yes

OpenStudy (anonymous):

I got 3/4

OpenStudy (amistre64):

\[2y\ y' + \left(3y+3x\ y'\right) (xy+1)^2 = 0\] \[2y\ y' + \left(3y+3x\ y'\right) ((xy)^2+2xy+1) = 0\] \[2y\ y' + \left(3y+3x\ y'\right) ((xy)^2+2xy+1) = 0\] \[2y\ y' + 3x^2y^3 +6xy^2+3y+(3x y' +6x^2y y'+3x^2y y') = 0\] \[3x^2y^3 +6xy^2+3y+(3x y' +6x^2y y'+3x^2y y'+2yy' ) = 0\] \[y'(3x +6x^2y+3x^2y+2y ) = -3x^2y^3 -6xy^2-3y\] \[y' = \frac{-3x^2y^3 -6xy^2-3y}{(3x +6x^2y+3x^2y+2y )};\ x=2,y=-1\] \[y' = \frac{-3(2)^2(-1)^3 -6(2)(-1)^2-3(-1)}{(3(2) +6(2)^2(-1)+3(2)^2(-1)+2(-1) )}\] \[y' = \frac{-3(4)(-1) -6(2)(1)+3}{(6 +6(4)(-1)+3(4)(-1)-2 )}\] \[y' = \frac{12 -12+3}{(6 -24-12-2 )}\] \[y' = \frac{3}{( -32 )}\] but i couldve missed it along the way